Haas Company manufactures and sells one product. The following information pertains to each of the companys first three yearComplete this question by entering your answers in the tabs below. Req 1 Reg 2A Req 2B Req ЗА Reg 3B Compute the companys brComplete this question by entering your answers in the tabs below. Reg 1 Req 2A Reg 2B Req ЗА Reg 3B Compute the unit productComplete this question by entering your answers in the tabs below. Reg 1 Reg 2A Reg 2B Req ЗА Reg 3B Prepare an income statemComplete this question by entering your answers in the tabs below. Reg 1 Reg 2A Reg 2B Req ЗА Req 3B Compute the unit productComplete this question by entering your answers in the tabs below. Req 1 Reg 2A Req 2B Reg 3A Reg 3B Prepare an income statem
Answer #1

Solution 1:

Contribution margin per unit = $64 - (29+21+5+1) = $8

Break even point = Fixed costs / Contribution margin per unit = (330000+150000) / 8 = 60,000 units

Solution 2:

a.

Computation of Unit Product Cost - Variable Costing
Particulars Year 1 Year 2 Year 3
Unit Product Cost:
Direct material $29.00 $29.00 $29.00
Direct Labor $21.00 $21.00 $21.00
Variable manufacturing overhead $5.00 $5.00 $5.00
Unit product cost $55.00 $55.00 $55.00

b.

Income Statement - Variable Costing - Haas Company
Particulars Per Unit Year 1 Year 2 Year 3
Sales $64 $38,40,000 $32,00,000 $41,60,000
Variable Costs:
Variable manufacturing Costs $55 $33,00,000 $27,50,000 $35,75,000
Variable Selling and administrative expenses $1 $60,000 $50,000 $65,000
Total Variable Costs $56 $33,60,000 $28,00,000 $36,40,000
Contribution Margin $8 $4,80,000 $4,00,000 $5,20,000
Fixed Expenses:
Fixed manufacturing overhead $3,30,000 $3,30,000 $3,30,000
Fixed Selling & Administrative Expenses $1,50,000 $1,50,000 $1,50,000
Net Operating Income $0 -$80,000 $40,000

Solution 3:

Computation of Unit Product Cost - Absorption Costing
Particulars Year 1 Year 2 Year 3
Unit Product Cost:
Direct material $29.00 $29.00 $29.00
Direct Labor $21.00 $21.00 $21.00
Variable manufacturing overhead $5.00 $5.00 $5.00
Fixed manufacturing overhead
($330,000 / Nos of unit produced)
$5.50 $4.40 $8.25
Unit Product Cost $60.50 $59.40 $63.25

b.

Income Statement - Absorption Costing - Haas Company
Particulars Year 1 Year 2 Year 3
Details Amount Details Amount Details Amount
Sales $38,40,000 $32,00,000 $41,60,000
Cost of Goods Sold:
Cost of goods produced $36,30,000 $44,55,000 $25,30,000
Less: Ending Inventory $0 $14,85,000 $0
Add: Opening Inventory $0 $36,30,000 $0 $29,70,000 $14,85,000 $40,15,000
Gross Profit $2,10,000 $2,30,000 $1,45,000
Variable Selling & Administrative Expenses $60,000 $50,000 $65,000
Fixed Selling & Administrative Expenses $1,50,000 $1,50,000 $1,50,000
Net Operating Income $0 $30,000 -$70,000